New symmetric designs for (79,27,9) Slavka Pfaff, Zagreb, Croatia

Keywords: symmetric design, automorphism group, orbit structure, fixed point (block) AMS Subject classification: 05B05 Abstract: Symmetric designs (79,27,9) with an automorphism group isomorphic to F_{39} × D_{8} are constructed and clasified.